Any chance you guys could help me with this is much appreciated!

Verify the identity?
[ ( cos()cot() ) / (1-sin()) ] - 1 = csc()
Thanks for the help, I believe I got all the correct parenthesis in place

Answer by stanbon(75887)

Verify the identity?
[ ( cos()cot() ) / (1-sin()) ] - 1 = csc()
----------
[cos(cos/sin)] / (1-sin) - 1 = 1/sin
--------------------------------------------
[cos^2 / sin] / (1-sin) - 1 = 1/sin
-----
Multiply both sides by sin to get:
---
cos^2 /(1-sin) - sin = 1
----
[(1-sin^2)/(1-sin) = 1+sin
-----
1+sin = 1+sin
==========================
Cheers,
Stan H.
------------------------